OpenStudy (anonymous):

Given f(x) = 5x - 6 and g(x) = x^2 - 5x + 6, find the function (f - g)(x). Use the ^ key for the exponent. Enter your answer as the following example: -2x^2+3x+10.

OpenStudy (anonymous):

your job is to compute \[(f-g)(x)=f(x)-g(x)=5x-6-(x^2-5x+6)\] remove the parentheses using the distributive property, then combine like terms is this clear?

OpenStudy (anonymous):

sorta

OpenStudy (anonymous):

just make sure to remove the parentheses carefully you should end up with \[5x-6-(x^2-5x+6)=5x-6-x^2+5x-6\] and then combine like terms

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

yw